How do you solve two-step equations? I tried answering 8 - 5w= -37 and got w=5.8 but when I checked it, it was wrong.
Answer by Alan3354(69443)

How do you solve two-step equations? I tried answering 8 - 5w= -37 and got w=5.8 but when I checked it, it was wrong.
--------------
8 - 5w= -37
Subtract 8
-5w = -45
5w = 45
w = 9